# Searching for the tetraneutron resonance on the lattice

### v1

@@ IV Summary
We have investigated the tetraneutron system confined in a cubic box using nuclear lattice effective field theory, with box sizes up to $L=30$ fm. Using the N3LO chiral EFT interaction, the confined $4n$ energy decreases gradually with increasing $L$ and remains concave over the explored volumes, showing no plateau that would be characteristic of a narrow resonance. For comparison, we also repeat the confined-spectrum calculation with an SU(4)-symmetric interaction fitted to the 1S0 phase shifts; the SU(4) results are consistent with the N3LO trend and are reported in 1.
To interpret possible $4n$ correlations in terms of two-dineutron dynamics, we employ the finite-volume dineutron approximation, where the lowest nnnn $S$-wave level in a periodic box is quasi-bound even though the infinite-volume $nn$ system corresponds to a virtual state. Within this framework, we validate the Lüscher-based extraction by studying $n$-$n$ and nn-$2n$ scattering. The ratio of scattering lengths $a_{n\text{-}2n}/a_{n\text{-}n}$ exhibits the expected universal behavior for near-threshold systems, with deviations only when $B_{2n}\approx 0.01$ MeV, where the proximity to the virtual-state regime amplifies finite-volume and lattice-spacing effects.
We then extract the 2n2n-$2n$ $S$-wave phase shifts from finite-volume $4n$ and $2n$ energies using Lüscher’s method together with the composite-dineutron finite-volume relation. At small momenta, the negative phase shifts indicate effective repulsion between dineutrons, consistent with effective field theory expectations near unitarity. In the intermediate window $p\simeq 55$$100$ MeV, $\delta(p)$ becomes positive and shows a shallow non-monotonic structure, with a maximum of order $10^{\circ}$ around $p\simeq 60$$84$ MeV, but without a rapid rise through $90^{\circ}$ typical of a narrow elastic resonance. This momentum band corresponds to box sizes $L\simeq 20$$15$ fm and confined $4n$ energies of 1.71.73.33.3 MeV, which lie close to the experimentally reported correlated $4n$ peak at $2.37\pm 0.38\text{(stat.)}\pm 0.44\text{(sys.)}$ MeV.
@@ V Acknowledgments
We are grateful for discussions with Kouichi Hagino, Hiroyuki Sagawa, Furong Xu, and the members of the NLEFT Collaboration. This work is supported by “the Fundamental Research Funds for the Central Universities”, National Natural Science Foundation of China under Grant No. U2541242 and 12435007. The work of SE is supported in part by the Scientific and Technological Research Council of Turkey (TUBITAK Project No. 123F464). The work of UGM was supported in by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC AdG EXOTIC, Grant Agreement No. 101018170), by the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2025PD0022). The work of YK was supported in part by the Institute for Basic Science (IBS-R031-D1). LW gratefully acknowledge the Computational resources provided by the HPC platform of Beihang University, National Supercomputing Center of Korea with supercomputing resources including technical support (KSC-2024-CHA-0001, KSC-2025-CHA-0004). SS and UGM gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC).
@@ V.1 Energy extrapolation
where $E_{\infty}^{(L)}$, $d^{(L)}$, and $c^{(L)}$ are fitting parameters. Details of the extrapolation procedure are provided in 1. We find that the dominant contribution to the interaction energy comes from the 1S0 channel, consistent with previous studies Lazauskas and Carbonell (2005). Other contributions, including the three-nucleon force, are negligibly smaller. For comparison, we also employ an SU(4) symmetric interaction Lu et al. (2019) fitted to the 1S0 phase shift, with the form:
@@ V.4 Phase shift
The neutron-neutron 1S0 scattering phase shift $\delta$, extracted from finite-volume two-neutron energies via Ls̈cher’s method, is shown in Fig. S5. Results obtained with the SU(4) and N3LO interactions are compared with experimental data Stoks et al. (1993).
@@
To test the universality of the dineutron approximation under pure contact interactions, we varied the dineutron binding energy from 0.1 MeV down to nearly zero at lattice spacings $a=1.32,\;1.64,\;1.97\;\text{fm}$, computed the relation between the dimensionless quantity $ap\cot\delta$ and $ap$ as shown in Fig. S6. Except for the system with a binding energy of 0.01 MeV, all results align with the expectations, indicating the validity of the dineutron approximation.
@@
Table S1: Scattering lengths and errors for different values of $a$ values ($B_{2n}=0.01$ MeV).
@@
value error value error value error
@@
Table S2: The ratio of neutron-dineutron to neutron-neutron scattering lengths for different lattice spacings and dineutron binding energies.
@@
$a_{n-2n}/a_{n-n}$ error $a_{n-2n}/a_{n-n}$ error $a_{n-2n}/a_{n-n}$ error
@@
Figure S7: Dineutron-dineutron $S$-wave scattering phase shift $\delta$ as a function of relative momentum $p$ extracted from finite-volume $4n$ and $2n$ energies using Lüscher’s method and the composite-dimer finite-volume relation in Eq. (6), for the SU(4) interaction. Different symbols correspond to the scanned near-threshold values of $B_{2n}$ (in MeV), and the shaded band indicates the momentum range associated with $L\simeq 20$$15$ fm.
We also calculated dineutron-dineutron $S$-wave scattering phase shifts for the SU(4) interaction, as shown in Fig. S7. In the momentum range 60-120 MeV, the phase shift also shows some attraction. Similar to the case with the N3LO interaction, no rapid rise of the phase shift to 90° is observed. The shaded region corresponds to $p\simeq$ 60–80 MeV, which for a box size of $L\simeq$ 20-15 fm yields a calculated tetraneutron energy range of $E\simeq 1.7$–3.0 MeV.

### v2

@@ IV Summary
We have investigated the tetraneutron system confined in a cubic box using nuclear lattice effective field theory, with box sizes up to $L=30$ fm. Using the N3LO chiral EFT interaction, the confined $4n$ energy decreases gradually with increasing $L$ and remains concave over the explored volumes, showing no plateau that would be characteristic of a narrow resonance. For comparison, we also repeat the confined-spectrum calculation with an SU(4)-symmetric interaction fitted to the 1S0 phase shifts; the SU(4) results are consistent with the N3LO trend and are reported in SM.
To interpret possible $4n$ correlations in terms of two-dineutron dynamics, we employ the finite-volume dineutron approximation, where the lowest nnnn $S$-wave level in a periodic box is quasibound even though the infinite-volume nnnn system corresponds to a virtual state. Within this framework, we validate the Lüscher-based extraction by studying nnnn and nn2n2n scattering. The ratio of scattering lengths $a_{n-2n}/a_{n-n}$ exhibits the expected universal behavior for near-threshold systems, with deviations only when $B_{2n}\approx 0.01$ MeV, where the proximity to the virtual-state regime amplifies finite-volume and lattice-spacing effects.
We then extract the $2n2n$ $S$-wave phase shifts from finite-volume $4n$ and $2n$ energies using Lüscher’s method together with the composite-dineutron finite-volume relation. At the smallest momenta, the extracted phase shifts show the largest sensitivity to $B_{2n}$. While for some of $B_{2n}$ values we observe slightly negative $\delta(p)$ in this regime, the low-momentum trend bends toward $\delta(p)\to 0$ as $B_{2n}\to 0$, and we therefore do not draw a firm conclusion from the sign of the lowest-momentum points. In the intermediate window $p\simeq 55$$100$ MeV, $\delta(p)$ becomes positive and shows a shallow nonmonotonic structure, with a maximum of order $10^{\circ}$ around $p\simeq 60$$84$ MeV, but without a rapid rise through $90^{\circ}$ typical of a narrow elastic resonance. This momentum band corresponds to box sizes $L\simeq 20$$15$ fm and confined $4n$ energies of 1.71.73.33.3 MeV, which lie close to the experimentally reported correlated $4n$ peak at $2.37\pm 0.38\text{(stat.)}\pm 0.44\text{(sys.)}$ MeV.
In the future, we plan to extend the NLEFT framework Elhatisari:2015iga to investigate the 8He($p,p^{4}$He)$4n$ reaction and understand the correlated structures observed. As a first step in this direction, NLEFT has recently been used to quantify four-neutron correlation patterns in 7H and 8He, motivated by quasifree knockout measurements Zhang:2025uin.
@@ V Acknowledgments
We are grateful for discussions with Kouichi Hagino, Hiroyuki Sagawa, Furong Xu, and the members of the NLEFT Collaboration. This work is supported by National Natural Science Foundation of China under Grant No. 12435007 and Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China. The work of S.E. is supported in part by the Scientific and Technological Research Council of Turkey (TUBITAK Project No. 123F464). The work of U.-G. M. was supported in by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (ERC AdG EXOTIC, Grant Agreement No. 101018170), by the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2025PD0022). The work of Y. K. was supported in part by the Institute for Basic Science (IBS-R031-D1). L.W. gratefully acknowledge the Computational resources provided by the HPC platform of Beihang University, the National Supercomputing Center of Korea with supercomputing resources including technical support (KSC-2024-CHA-0001, KSC-2025-CHA-0004). $S.$ and U.-G. M. gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this project by providing computing time on the GCS Supercomputer JUWELS at the Jülich Supercomputing Centre (JSC).
Data availability—The data that support the findings of this article are openly available Wu2026.
@@ V.1 Energy extrapolation
where $E_{\infty}^{(L)}$, $d^{(L)}$, and $c^{(L)}$ are fitting parameters. We find that the dominant contribution to the interaction energy comes from the 1S0 channel, consistent with previous studies Lazauskas and Carbonell (2005). Other contributions, including the three-nucleon force, are negligibly smaller. For comparison, we also employ an SU(4) symmetric interaction Lu et al. (2019) fitted to the 1S0 phase shift, with the form:
@@ V.4 Phase shift
The neutron-neutron 1S0 scattering phase shift $\delta$, extracted from finite-volume two-neutron energies via Lüscher’s method, is shown in Fig. S5. Results obtained with the SU(4) and N3LO interactions are compared with experimental data Stoks et al. (1993).
@@
To test the universality of the dineutron approximation under pure contact interactions, we varied the dineutron binding energy from 0.1 MeV down to nearly zero at lattice spacings $a=1.32,\;1.64,\;1.97\;\text{fm}$, computed the relation between the dimensionless quantity $ap\cot\delta$ and apap as shown in Fig. S6. Except for the system with a binding energy of 0.01 MeV, all results align with the expectations, indicating the validity of the dineutron approximation.
@@
Table S1: Scattering lengths for different values of $a$ values ($B_{2n}=0.01$ MeV). The column of uncertainty indicates the statistical uncertainty from the fits to the effective-range expansion.
@@
value uncertainty value uncertainty value uncertainty
@@
Table S2: The ratio of neutron-dineutron to neutron-neutron scattering lengths for different lattice spacings and dineutron binding energies. The column of uncertainty indicates the statistical uncertainty from the fits to the effective-range expansion.
@@
$a_{n-2n}/a_{n-n}$ uncertainty $a_{n-2n}/a_{n-n}$ uncertainty $a_{n-2n}/a_{n-n}$ uncertainty
@@
Figure S7: Dineutron-dineutron $S$-wave scattering phase shift $\delta$ as a function of relative momentum $p$ extracted from finite-volume $4n$ and $2n$ energies using Lüscher’s method and the composite-dimer finite-volume relation for the SU(4) interaction. Different symbols correspond to the scanned near-threshold values of $B_{2n}$ (in MeV), and the shaded band indicates the momentum range associated with $L\simeq 20$$15$ fm.
We also calculated dineutron-dineutron $S$-wave scattering phase shifts for the SU(4) interaction, as shown in Fig. S7. In the momentum range 60-120 MeV, the phase shift also shows some attraction. Similar to the case with the N3LO interaction, no rapid rise of the phase shift to $90^{\circ}$ is observed. The shaded region corresponds to $p\simeq$ 60–-80 MeV, which for a box size of $L\simeq$ 20–15 fm yields a calculated tetraneutron energy range of $E\simeq 1.7$–3.0 MeV.